1. A person has 1000 J of kinetic energy. What is their velocity if they have a mass of 100kg?
1
Expert's answer
2021-05-03T17:27:52-0400

We can find the velocity of the person from the definition of the kinetic energy:


"KE=\dfrac{1}{2}mv^2,""v=\sqrt{\dfrac{2KE}{m}},""v=\sqrt{\dfrac{2\cdot1000\ J}{100\ kg}}=4.47\ \dfrac{m}{s}."